OpenStudy (anonymous):

Write the explicit formula for the geometric sequence. a1 = -5 a2 = 20 a3 = -80

OpenStudy (anonymous):

It looks like each successive term is the previous one multiplied by -4, so that's the common ratio between terms. What do you know about the general form of a geometric sequence?

OpenStudy (anonymous):

uhm nothing rellee

OpenStudy (anonymous):

maybe you could just give me the answer?

OpenStudy (anonymous):

A geometric sequence is of the form \[\{ar^n\}\] where r is the common ratio and n starts with, say, 0 and goes up by one for each successive term. For example, the geometric sequence \[\left\{2\left(\frac{1}{2}\right)^n\right\} = \left\{ 2, 1, \frac{1}{2}, \frac{1}{4},\ldots \right\}\] So, when n = 0, you have 2, n = 1, you get 1, and so on. (Each successive term is divided by 2, or multiplied by 1/2.) For your sequence, clearly the first term (when n=0) is -5, and r = -4. Think you can write up a general sequence using that info?

OpenStudy (anonymous):

Giving you the answer would hardly help you understand what you're doing.

OpenStudy (anonymous):

an = -5 • (-4)n ?

OpenStudy (anonymous):

That's right. Just remember the n is an exponent.

OpenStudy (anonymous):

ok wait i have another question tho!

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

Write the recursive formula for the geometric sequence. a1 = -2 a2 = 8 a3 = -32

OpenStudy (anonymous):

Following the same process, I'd first write the sequence as a regular geometric sequence, which you should be able to understand is \[a_n = -2 (-4)^n,\] right?

OpenStudy (anonymous):

riighhhht..

OpenStudy (anonymous):

That means that \[a_{n+1} = -2(-4)^{n+1}\] If you're familiar with properties of exponents, you should be able to easily express a_(n+1) in terms of a_n.

OpenStudy (anonymous):

nopee

OpenStudy (anonymous):

\[a_{n+1} = -2(-4)^{n+1} = -2(-4)^n \cdot (-4)^1\] Look familiar?

OpenStudy (anonymous):

honestly noo...

OpenStudy (anonymous):

Well, in any case, can you see that you can substitute a_n into that last equation? I gave you a_n earlier, so you shouldn't have any problems with finding it.